OpenStudy (anonymous):

I have absolutly no clue on this one o-o Can you help please :o @amistre64 @acxbox22

OpenStudy (amistre64):

its going to help to have a clue ...

OpenStudy (amistre64):

i simply dont get paid enough to do all the thinking :)

OpenStudy (anonymous):

Im sorry but I have no idea what that means xD

OpenStudy (amistre64):

is the sequence arithmetic or geometric? how do we tell?

OpenStudy (anonymous):

geometric

OpenStudy (amistre64):

good, and our common ratio is?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

hmm 3r = 12, 12r = 48 what is r?

OpenStudy (anonymous):

R is ratio right? So its being multiplied by 3

OpenStudy (amistre64):

3*3 = 9, not 12 12*3 = 36, not 48 lets say r=4 the formula is:\[S_n=a_1\frac{1-r^n}{1-r}\]

OpenStudy (amistre64):

all we really need to determine now is the value of n

OpenStudy (anonymous):

TIMES 4 IM SORRY I READ MY CALCULATOR WRONG Its 4 i know what 3*3 is xDD

OpenStudy (amistre64):

we know the first term, we know r, and we know the value of the last term giving us: 49152 = 3(4)^(n-1) and we solve for n

OpenStudy (amistre64):

preethas here so we might be getting filmed :)

OpenStudy (anonymous):

...filmed? For what?

OpenStudy (acxbox22):

lol

OpenStudy (anonymous):

and hold on lemme work out that equation

OpenStudy (acxbox22):

i wanta be part of da action!

OpenStudy (amistre64):

when im older im going to write my memoirs ... this will be proof that i failed in life lol

OpenStudy (acxbox22):

LOLOLOLOL

OpenStudy (anonymous):

idk what to say to that xD And I have 49152^n-1 but idk what n is D:

OpenStudy (amistre64):

49152 = 3(4)^(n-1) 49152/3 = 4^n 4^(-1) 4*49152/3 = 4^n log(4*49152/3) = n log(4) log(4*49152/3) / log(4) = n

OpenStudy (anonymous):

woah what no I have 49152=12^n-1

OpenStudy (anonymous):

AMI I AM SO LOST. THAT IS SO MANY NUMBERS

OpenStudy (amistre64):

i used the acceptable amount of numbers, even tried to put some extra ones in there to boot

OpenStudy (amistre64):

does 3(4)^7 = whatever that last number was? 49152?

OpenStudy (anonymous):

yes it does

OpenStudy (amistre64):

lets assume im right :) now we work the formula \[S_8=\frac{3(1-4^8)}{1-4}\]

OpenStudy (anonymous):

Hi mom!

OpenStudy (amistre64):

how many times do i have to tell you!! im not your mother!!!

OpenStudy (anonymous):

21842

OpenStudy (amistre64):

you might want to change the batteries in your calculator

OpenStudy (amistre64):

3(1-4^8)/(1-4) http://www.wolframalpha.com/input/?i=3%281-4%5E8%29%2F%281-4%29

OpenStudy (anonymous):

-196609 o.O

OpenStudy (amistre64):

it begins with an A and ends with a 5 ....

OpenStudy (anonymous):

65 thousand is a lot different from whatever the hell my number was o-o

OpenStudy (amistre64):

i agree, hence the reference to a change of batteries :)

OpenStudy (anonymous):

xD Okay thank you ami ^_^

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

wait really quick what was the formula for this?

OpenStudy (acxbox22):

go to the 10th post down :)

OpenStudy (anonymous):

kk got it ty ray ;p

OpenStudy (anonymous):

and ami :o <33